OpenStudy (anonymous):

jack mows his yard in 4hr. sue mows in 3hr. how long will it take both of them to mow the same yard? 2hr?

OpenStudy (slaaibak):

IIRC, it's something like 1/3 + 1/4 = 1/(new rate) = 7/12 Therefore, answer is 12/7 hours. Not sure tho :/

OpenStudy (anonymous):

i dont think my answer is correct im still working on it, its confusing

OpenStudy (slaaibak):

I think my answer is, it makes logically sense. It would definately be less than 2 hours, since sue mows faster than jack. similarly, it would definately be less than 1.5, since jack mows longer than 2 hours. \[1.5 < {12 \over 7} < 2\] Therefore my answer and method is correct. It takes them both 12/7 hours.

OpenStudy (slaaibak):

*similarly, it would definately be MORE than 1.5, since jack mows longer than sue. Sorry,typo*

hero (hero):

\[\frac{a \times b}{a + b}\]

hero (hero):

\[\frac{4 \times 3}{4 + 3} = \frac{12}{7}\]